Fermi National Laboratory
 

Operation Summary for September 3, 2010 - September 6, 2010

 

The Integrated Luminosity for the period from 8/30/10 to 9/6/10 was 37.14 pb-11.
NuMI reported receiving 6.52E18 protons on target during this same period.

Friday Sept 3

The day shift began with the Tevatron (TeV) F-sector in access, with Operations monitoring the Recycler stash (283.2E10) and the Antiproton Source (Pbar) stack (12.6E10), with the Linac and Booster conducting maintenance work, with the Main Injector (MI), NuMI, and MiniBooNE off due to the F-sector access and Booster work, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, vacuum technicians worked all shift to fix the TeV vacuum leak. Booster, NuMI, and Pbar took advantage of the downtime to access and perform maintenance.

During the evening shift, Operations established beam to MiniBooNE at 4:36 PM and to NuMI and Pbar by 4:50 PM. Operators had the TeV on and in a dry squeeze at 10:48 PM. TeV personnel conducted orbit studies.

Saturday Sept 4

The midnight shift began with Operations monitoring the stash (299.6E10), with Pbar stacking (104.E10), and with MiniBooNE taking beam.

During the midnight shift, TeV orbit studies concluded and operations put the TeV into HEP shot setup. Operations established store 8055 at 3:21 AM with an average initial luminosity (AIL) of 198.2E30.

During the day shift, a Linac RF station (LRF1) tripped off at 9:52 AM and required Linac personnel to come in and investigate; they had the station on by 11:43 AM. Operations terminated store 8055 at 1:32 PM, put the TeV into shot setup, and established store 8056 at 2:51 PM with an AIL of 246.1E30.

During the evening shift, A MI RF station (MIRF2) tripped off five times; it reset. RF technicians replaced a bias power supply for MIRF6 and reported that MIRF2 was tripping because MIRF6 was down.

Sunday Sept 5

The midnight shift began with Operations monitoring store 8056 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that BRF6 had tripped off at 2:20 AM. An operator called in a RF technician person who discovered that its yard breaker had tripped off; the Duty Electrician had it reset by 3:49 AM. Operations terminated store 8056 at 5:19 AM, put the TeV into shot setup, and established store 8057 at 6:48 AM with an AIL of 253E30.

During the day shift, at 9:55 AM, MIRF6 tripped off and would not reset; operators bypassed the station.

During the evening shift, Operations terminated store 8057 at 8:33 PM, put the TeV into shot setup, and established store 8058 at 10:49 PM with an AIL of 261.7E30.

Monday Sept 6

The midnight shift began with Operations monitoring store 8058 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported a 3E30 drop of luminosity in store 8058 at 1:18 AM, which corresponded with a problem on TeV RF station (TRF6).

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/07/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory